OpenStudy (anonymous):

logarithmic expressions: solve to the nearest thousandth.. \[5^{1 + x} = 2^{1-x}\]

OpenStudy (anonymous):

start with \[(x+1)\ln(5)=(1-x)\ln(2)\] and then do a bunch of algebra recalling that \(\ln(5)\) and \(\ln(2)\) are just some constants

OpenStudy (anonymous):

i can walk you through it if you like, but it is really algebra to get \(x\) by itself on one side of the equal sign

OpenStudy (anonymous):

yes please, cause i got through first two equations of the question its just this one i have left.

OpenStudy (anonymous):

\[(x+1)\ln(5)=(1-x)\ln(2)\] \[x\ln(5)+\ln(5)=\ln(2)-x\ln(2)\] \[x\ln(5)+x\ln(2)=\ln(2)-\ln(5)\] \[x(\ln(5)+\ln(2))=\ln(2)-\ln(5)\] \[x=\frac{\ln(2)-\ln(5)}{\ln(5)+\ln(2)}\]

OpenStudy (anonymous):

then a calculator, because i have no idea what this is

OpenStudy (anonymous):

LOL okay. wait but if you bring (ln(5)+ln(2)) to the right side doesn't it become (ln(5)-ln(2)) ??

OpenStudy (anonymous):

oh... wait... nvm!

OpenStudy (anonymous):

not sure what you are asking steps are 1) distributive property 2) add \(x\ln(2)\) to both sides 3) subtract \(\ln(5)\) from both sides 4) factor the \(x\) 5) divide

OpenStudy (anonymous):

lol sorry i confused myself for a second there.

OpenStudy (anonymous):

easy to do but thing of \(\ln(5)=a\) and \(\ln(2)=b\) they are just constants

OpenStudy (anonymous):

*think